A chemist has two salt solutions. One solution is 20% salt and the other if 50% salt. How many liters of each must be used to produce 600 liters of a solution that is 10% salt?
Answer by checkley71(8403)

YOU CAN'T MIX A 20% SOLUTION WITH A 50% SOLUTION AND GET A 10% SOLUTION WITHOUT ADDING WATER OR SOME OTHER NON-SALT SOLUTION. THE HIGH & LOW SOLUTIONS ARE 599*.2+1*.5=600X OR 119.8+.5=600X OR 120.3=600X OR 120.3/600=X OR X=20.09% SOLUTION AND A MAX SOLUTION OF 599*.5+1*.2=600X OR 299.5+.2=600X OR 299.7/600=X OR X=49.95%.